2. S.S. Corporation’s bonds will mature in 15 years. The bonds have a face value of $1,000 and

an 6.5 percent coupon rate, paid semiannually. The price of the bonds is $1,050. What is the yield to maturity?

5.99% Please show all worked typed. I can not see much on mobile
Answer #1

2. The face value of the bond is $1000,

the bond is being sold currently at a price of $1050, the coupon rate is 6.5%,

As the bond is compounded semi-annually,

FV = $1000

PV = ($1050)

N = 30 YEARS

PMT = $32.5

INPUTTING THESE IN THE FINANCIAL CALCULATOR WE GET,

I/Y AS = 2.99%

ANNUAL RATE OF INTEREST IS = 5.99%,

ALTERNATIVELY,

32.5 /(1 +Y)^1 + 32.5/(1+Y)^2 +..... 32.5/(1+Y)30 = $1050

SOLVING WHICH WE GET,

I/Y = 5.99% when compounded annually.